Determine a,b,c of the quadratic equation 2x²4x=8
1
Expert's answer
2021-10-25T03:14:04-0400

Let us determine "a,b,c" of the quadratic equation "2x\u00b2+4x=8." The equation "2x\u00b2+4x=8" is equivalent to "2x\u00b2+4x-8=0," and hence "a=2, b=4, c=-8."